### v3

# Preprint Manuscript

Preprint Manuscript
Third Version
CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM OVER $\mathbb{N}$ STARTS AT $\delta=3$
Milan Rosko
(Oktober 2025)
We prove that Hilbert’s Tenth Problem over $\mathbb{N}$ remains undecidable when restricted to cubic equations (degree $\leq 3$), resolving the open case $\delta=3$ identified by Jones (1982) and establishing sharpness against the decidability barrier at $\delta=2$ (Lagrange’s four-square theorem). For any consistent, recursively axiomatizable theory $T$ with Gödel sentence $G_{T}$, we effectively construct a single polynomial $P(x_{1},\ldots,x_{m})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that $T\vdash G_{T}$ if and only if $\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0$. Our reduction proceeds through four stages with explicit degree and variable accounting. Proof-sequence encoding via Diophantine $\beta$-function and Zeckendorf representation yields $O(KN)$ quadratic constraints ($K=O(\log(\max_{i}f_{i}))$, NN = proof length). Guard-gadgets wrap each constraint $E(\mathbf{x})=0$ as $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$, preserving degree $\leq 3$ with $O(KN^{3})$ variables. Sum-of-squares aggregation $P_{\text{merged}}=\sum_{i}P_{i}^{2}$ produces degree $\leq 6$ with $O(KN^{3})$ monomials. Recursive monomial shielding reduces degree to $\leq 3$ in $O(\log d)$ rounds, adding $O(K^{3}N^{3})$ variables total.
Keywords: Proof Theory, Logic, Hilbert’s Tenth Problem, Zeckendorf’s Theorem
Open Problem, Constructive Mathematics, Fibonacci, Cubic equations
Primary MSC: 11U05, 03F40, 03D35
Secondary MSC: 11D72, 03B25, 11Y16
Milan Rosko is from University of Hagen, Germany
Email: Q1012878 @@ studium.fernuni-hagen.de – ORCID: 0009-0003-1363-7158
Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
1 Corrigendum
1.1 First Corrigendum
The initial version encoded a Gödel sentence as a Diophantine system rather than a single equation. The ambition to present a concrete standalone equation introduced an unwarranted bias into the construction. Subsequent algorithmic approaches yielded a massive closed form, but the reduction diverted attention from the central epistemic questions.
1.2 Second Corrigendum
In the second revision, we pivoted to explicit witness construction, ensuring all constraints remain genuinely cubic, and grounded the argument in logical foundations rather than forcing a premature encoding.
1.3 Third Corrigendum
In this third—still preliminary—version, we resolve ambiguities between logical and mathematical layers ($\mathbb{Z}$ versus $\mathbb{N}$), and improve on bookkeeping. The first-order argument is developed independently of its Diophantine realization. We provide a high-level overview at the outset, then verify each layer on its own terms.
[…]
We now adopt a direct approach: all variables, constraints, and solutions are understood to inhabit $\mathbb{N}$ from the outset. The $\beta$-function (Theorem 3.1), four-squares encoding (Lemma 4.2), and Zeckendorf representation (Definition 3.3) operate naturally over $\mathbb{N}$. The guard-gadget framework (Definition 3.6) enforces positivity via $u=1+v^{2}\geq 1$ without recourse to negative integers. This interpretation aligns with the historical formulation of Hilbert’s Tenth Problem over $\mathbb{N}$ and avoids unnecessary detours through $\mathbb{Z}$. The approach is straightforward: Turing machines operate with discrete symbols.
1.4 Context
This paper shares ideas with different work [^1] The Fractal Logic of $\Phi$-adic Recursion, arXiv:2510.08934, [2025a], in that a compression-undecidability relationship is visible in the context of extensions from $\mathbb{N}$ to $\mathbb{Z}$. The additional compression offered by negative numbers permits lower polynomial degrees in Diophantine equations, but at the cost of increased uncertainty between powers. The equation $x^{2}=4$ is uniquely solved in $\mathbb{N}$ by $x=2$, while in $\mathbb{Z}$ the solution set comprises $\{-2,2\}$. More dramatically, $x^{4}=16$ admits four solutions $\{\pm 2,\pm 2i\}$ in $\mathbb{C}$, with the real solutions $\pm 2$ already exhibiting the doubling. Future work will focus on this: The ambiguity is not a deficiency but a necessary consequence of Kolmogorov [1968] complexity: syntactic reduction of degrees is purchased through semantic branching. Results possessing nontrivial structure in $\mathbb{N}$ collapse in $\mathbb{Z}$ to trivial symmetries, whereby the displaced complexity migrates into sign ambiguity.
2 Overview
2.1 The Problem
Previously [^2] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, arXiv:2509.10382, [2025b], we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability. We now apply this method to produce undecidable Diophantine equations of degree three, thereby advancing “Hilbert’s Tenth Problem” [^3] As posed by Hilbert [1900]. into the cubic regime ($\delta=3$) and resolving a previously open case: We cite from the article “Universal Diophantine Equation”, by Jones [1982]:
We know that there exist universal diophantine equations of degree 4 […] but none of degree 2. […] Thus the only open problem as regards the degree is $\delta$ = 3. We do not know whether Hilbert’s tenth problem is decidable or undecidable for equations of degree 3.
Hence, we define the open question:
Problem 2.1 (Cubic Remainder of Hilbert’s 10th Problem).
Let $\delta=3$ denote the set of all polynomial Diophantine equations of the form
$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{N}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (2.1)
Question: Is there an algorithm that, given $P$ and $\delta$ = 3, decides whether
$\exists(x_{1},\ldots,x_{n})\in\mathbb{N}^{n}:P(x_{1},\ldots,x_{n})=0?$ (2.2)
2.2 Roadmap
[…]
The restriction to $\mathbb{N}$ rather than $\mathbb{Z}$ is essential. Cubic equations over $\mathbb{Z}$ admit decidability through Thue equations (Baker’s transcendence methods), Hasse principle obstructions, and elliptic curve descent via Mordell–Weil theory. Undecidability over $\mathbb{N}$ exploits combinatorial complexity introduced by non-negativity constraints, enforced polynomially through guard-gadgets: replacing constraints $E(\mathbf{x})=0$ of degree at most two with the guarded system $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$ maintains cubic degree while forcing $u\geq 1$. Naive activation variables or difference-of-squares encodings produce degree-four terms, necessitating this construction.
Stage Configuration Implication
Step 1 $\mathcal{A}_{1}$ exists $\delta=2$ decidable (four-squares)
Step 2 $\mathcal{A}_{1}\land\mathcal{A}_{2}$ exist Can reduce any $\delta=6$ to equivalent $\delta=3$
Step 3a Assume $\mathcal{A}_{3}$ exists Contradiction: $\mathcal{A}_{1}\land\mathcal{A}_{2}\land\mathcal{A}_{3}$ decides PA
Step 3b Reject hypothesis Conclusion: $\delta=3$ undecidable over $\mathbb{N}$
Table 1: Logical structure of the impossibility proof.
The technical realization is a witness of a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds Zeckendorf digit count. Every intermediate constraint—after guard-wrapping—has degree $\leq 3$ (by explicit bookkeeping, see 2); sum-of-squares merging yields degree $\leq 6$, reduced to cubic via monomial shielding (Definition 6.1) with $O(K^{3}N^{3})$ auxiliary variables
$P_{\text{merged}}=\sum_{i}P_{i}^{2}\in\delta(6)$ (2.3)
followed by recursive monomial shielding: each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ factors as $p\cdot q$ with $\delta(p),\delta(q)\leq 3$, introducing auxiliary $w$ via $w=p\cdot q$. Iteration terminates in $O(1)$ rounds, yielding a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $O(K^{3}N^{3})$ variables. The variable inflation from systems to single equations reflects the cost of degree-preserving aggregation.
Our proof by contradiction states: No universal cubic equation $P_{\text{univ}}(n,\mathbf{x})$ can encode all halting problems: such a polynomial would render Chaitin [1975] constant
$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.4)
computable by enumeration. Our construction avoids this through non-uniformity: $P_{\text{hard}}$ depends on the uncomputable proof length $N$. This applies equally to systems and single equations, though systems admit sparse structure while single equations become dense under sum-of-squares.
[…]
The cubic barrier represents a genuine phase transition in Diophantine complexity over $\mathbb{N}$, arising from arithmetic structure, non-negativity, and computational irreducibility: systems achieve cubic degree through encoding design, while single equations require degree-reduction post-processing.
Figure 1: A high-level overview of the reduction from provability in Peano Arithmetic to the solvability of cubic Diophantine systems. The construction proceeds through seven stages, each maintaining polynomial bounds on variable count and equation degree.
2.2.1 Encoding Proof Sequences (Stages 1–3)
A proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ is encoded via Gödel numbers $f_{i}\in\mathbb{N}$. The sequence is stored using two complementary representations:
  1. 1.
    $\beta$-Function Storage (Theorem 3.1): Existentially quantify over parameters $c,d\in\mathbb{N}$ satisfying
    $\beta(c,d,i)=f_{i}\quad\text{for all }i\in\{1,\ldots,N\},$ (2.5)
    where $\beta(c,d,i)$ denotes the remainder when dividing $c$ by $1+(i+1)d$. This is enforced via quadratic constraints (Lemma 4.1):
    $c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (2.6)
[…]
  2. 2.
    Zeckendorf Digits (Definition 3.3): Each $f_{i}$ is represented as a sum of non-consecutive Fibonacci numbers:
    $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (2.8)
    This representation is unique and enables arithmetic verification without carry propagation (Exposition Exposition).
2.2.2 Proof Verification (Stage 4)
Each line $i$ must be justified either as an axiom or by modus ponens from earlier lines. This disjunction is encoded via boolean activation variables:
  1. 1.
    Axiom Test (Lemma 4.5): For a finite axiom set $\{g_{1},\ldots,g_{M}\}$, the constraint
    $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0,\quad b_{ax,i}\in\{0,1\}$ (2.9)
    forces $f_{i}\in\{g_{1},\ldots,g_{M}\}$ when $b_{ax,i}=1$ (since a sum of squares vanishes iff each term vanishes).
  2. 2.
    Modus Ponens Test (Lemma 4.7): For each triple $(i,j,k)$ with $j,k<i$, the constraint
    $b_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0,\quad b_{mp,i,j,k}\in\{0,1\}$ (2.10)
    forces $f_{i}=f_{j}+f_{k}$ when $b_{mp,i,j,k}=1$. The Zeckendorf digits of $f_{i}$ automatically satisfy the non-adjacency constraint by uniqueness (Remark Exposition).
  3. 3.
    Exactly-One Justification (Lemma 4.11): The constraint
    $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1$ (2.11)
    ensures that exactly one justification mechanism activates per line.
2.2.3 Degree Management (Stages 5–7)
The proof verification constraints (Stage 4) produce equations of degree $\leq 3$, but merging them into a single equation via sum-of-squares yields degree $6$. We employ two techniques:
  1. 1.
    Guard-Gadgets (Definition 3.6): Each constraint $E(\mathbf{x})=0$ with $\delta(E)\leq 2$ is replaced by the guarded system
    $u\cdot E(\mathbf{x})$ $=0,$ (2.12)
    $u-1-v^{2}$ $=0,$ (2.13)
    where $u,v\in\mathbb{N}$. The second equation forces $u=1+v^{2}\geq 1$, hence $u\neq 0$, so the first equation implies $E(\mathbf{x})=0$. This increases degree to at most $\max(1+2,2)=3$.
  2. 2.
    Jones Factorization (Theorem 5.1): After merging via sum-of-squares (yielding $P_{\text{merged}}$ of degree $6$), we apply recursive monomial shielding. Each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ is factored as $p\cdot q$ where $\delta(p),\delta(q)\leq 3$, introducing auxiliary variables $w$ via constraints $w=p\cdot q$ of degree $\leq 3$. This process (Algorithm 6.3) introduces $O(m\log d)$ variables for a polynomial with $m$ monomials of degree $\leq d$.
2.3 Main Result
Theorem 2.2 (Restatement of Theorem 5.2).
Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be its Gödel sentence. There exists a polynomial $P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that:
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (2.14)
The polynomial has $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds the Fibonacci representation size.
2.4 On Single Cubic Equations
While our construction produces a system of cubic equations, one might ask whether a single cubic equation could suffice. The following obstruction shows this is impossible.
Proposition 2.3 (Impossibility of Universal Cubic Equation).
There exists no polynomial $P_{\text{univ}}(n,\mathbf{x})\in\mathbb{Z}[n,\mathbf{x}]$ within an $\omega$-consistent formal system of degree $\leq 3$ such that for all Turing machines $M$ (encoded as $n\in\mathbb{N}$):
$M\text{ halts}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{k}:P_{\text{univ}}(n,\mathbf{x})=0.$ (2.15)
Proof sketch.
Assume such $P_{\text{univ}}$ exists. For each machine $M$ encoded as $n$, enumerate solutions $\mathbf{x}\in\mathbb{N}^{k}$ in order of increasing norm $|\mathbf{x}|_{1}=\sum_{i}x_{i}$. By the degree-$3$ constraint, the polynomial $P_{\text{univ}}(n,\mathbf{x})$ has at most $O(k^{3})$ monomials, so checking whether $P_{\text{univ}}(n,\mathbf{x})=0$ takes polynomial time in $|\mathbf{x}|_{1}$.
For halting machines, the smallest solution $\mathbf{x}^{*}$ satisfies $|\mathbf{x}^{*}|_{1}=O(\text{poly}(n))$ (by the polynomial encoding of bounded computations). Thus, we could compute Chaitin [1975] constant.
$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.16)
by enumerating solutions up to polynomial size—contradicting the incomputability of $\Omega$, hence no such $P_{\text{univ}}$ exists. ∎
Remark.
Our construction avoids this obstruction by producing a machine-dependent system $\Phi_{M}(\mathbf{x})$. The proof length $N$ (and hence the variable count $m=O(K^{3}N^{3})$) is not computable from the machine description alone—this non-uniformity is essential. The system encodes "there exists a proof of length NN", not "the shortest proof has length NN".
2.5 Prior Work and Roadmap
We recall:
  1. 1.
    Robinson et al. [1961]: Showed exponential Diophantine relations are listable, reducing Hilbert’s 10th to finding a Diophantine representation of exponentiation. While their construction used higher degrees, our guard-gadget technique demonstrates degree $\leq 3$ suffices for universal representation over $\mathbb{N}$.
  2. 2.
    Jones [1982]: Achieved degree $\leq 4$ via direct MRDP reduction. Our guard-gadget technique (Definition 3.6) reduces this to degree $\leq 3$ by systematically shielding quadratic constraints.
  3. 3.
    Matiyasevich [1970]: Proved undecidability for degree $\leq 4$ (settling the “vanilla” Hilbert’s 10th Problem). Our result shows the degree barrier can be lowered by one, addressing a natural question raised by Jones [1982].
  4. 4.
    Hilbert’s Tenth Problem and its variants for rings, number fields, and related structures remain central questions in mathematics [Alpöge et al., 2025]. However, from a computational perspective, structures like $\mathbb{Z}$ reduce to $\mathbb{N}$ via standard encodings. Mathematical results, see [Poonen, 2009], concern definability complexity in strong foundational theories, which is orthogonal to our focus on bounded provability in weak arithmetics.
The remainder of the paper fills in the technical details:
  1. 1.
    Section 3 (Preliminaries): Establishes encoding tools ($\beta$-function, four-squares, Zeckendorf, guard-gadgets).
  2. 2.
    Section 4 (Construction): Builds the cubic system via proof sequence encoding, axiom testing, and modus ponens verification.
  3. 3.
    Section 5 (Aggregation): Merges the system via sum-of-squares and applies Jones factorization.
  4. 4.
    Section 6 (Computation): Provides algorithmic details for degree reduction.
  5. 5.
    Section 7 (Result): Proves soundness and completeness bidirectionally.
Each stage is constructed to be effective (algorithmically computable), polynomial-time (in the proof parameters), and degree-preserving (maintaining $\delta\leq 3$ after Stage 5).
3 Preliminaries
3.1 The Gödel Beta-Function
The following logic allows encoding arbitrary finite sequences as Diophantine constraints:
Procedure 3.1 (Diophantine $\beta$-Function).
By the work of Gödel [1931] (using the “Chinese remainder”) and as exposited in Matiyasevich [1993], there exists a polynomial $\beta(c,d,i)$ over $\mathbb{N}$ such that for any finite sequence $a_{0},a_{1},\ldots,a_{n}\in\mathbb{N}$, there exist $c,d\in\mathbb{N}$ with
$\beta(c,d,i)=a_{i}\quad\text{for all }i\in\{0,1,\ldots,n\}.$ (3.17)
Explicitly, $\beta(c,d,i)=r_{i}$, where $r_{i}$ is the remainder and $q_{i}\in\mathbb{N}$ is the quotient when dividing $c$ by $(1+(i+1)d)$:
$c=q_{i}\cdot(1+(i+1)d)+r_{i},\quad 0\leq r_{i}<1+(i+1)d.$ (3.18)
Remark.
The “Chinese remainder” allows us to choose $d$ such that the moduli $\{1+(i+1)d:0\leq i\leq n\}$ are pairwise coprime. A suitable choice is
$d=\text{lcm}(1,2,\ldots,n+1),$ (3.19)
which ensures that for any $i\neq j$, the moduli $1+(i+1)d$ and $1+(j+1)d$ differ by (ji)d(j-i)d, and this difference shares no common factor with either modulus, since
$1+(k+1)d\equiv 1\pmod{p}\quad\text{for all primes}\quad p\mid d\quad\text{with}\quad p\leq n+1.$
Hence we construct $c$ via CRT to satisfy $c\equiv a_{i}\pmod{1+(i+1)d}$ for all $i\leq n$. To verify pairwise coprimality with $d=\text{lcm}(1,\ldots,n+1)$: for $i<j\leq n$, we have
$\gcd(1+(i+1)d,1+(j+1)d)\mid\gcd(1+(i+1)d,(j-i)d).$ (3.20)
Since $(j-i)\leq n$ and $d$ is divisible by all integers up to $n+1$, while $1+(i+1)d\equiv 1\pmod{p}$ for any prime $p\mid d$ with $p\leq n+1$, the greatest common divisor equals $1$. This ensures the CRT applies. To express $\beta(c,d,i)=f_{i}$ as polynomial constraints, we introduce variables $r_{i},q_{i}$ and enforce:
$c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (3.21)
$0$ $\leq r_{i}<1+(i+1)d.$ (3.22)
The second challenge, after reducing $\Sigma_{1}$ logic to $\Delta_{0}$ [^4] Foundational reference by Kleene [1943] introducing concepts underlying the arithmetic hierarchy, is encoding inequality polynomially. A naive preliminary approach using $r_{i}=s_{i}^{2}$ to force $r_{i}\geq 0$ failed because: (i) the $\beta$-construction produces remainders $r_{i}=c\bmod(1+(i+1)d)$ that need not be perfect squares (e.g., $r_{i}=2$ is not a square), and (ii) encoding the upper bound $r_{i}<1+(i+1)d$ via $(1+(i+1)d-r_{i}-1)=t_{i}^{2}$ would impose the same impossible requirement. The “four-squares” encoding resolves both issues simultaneously without requiring any specific number to be a perfect square.
Lemma 3.2 (Four-Squares Bound Encoding).
The constraint $0\leq r<m$ over $\mathbb{N}$ is equivalent to the existence of integers $a,b,c,e,\alpha,\beta,\gamma,\delta\in\mathbb{N}$ satisfying:
$r$ $=a^{2}+b^{2}+c^{2}+e^{2},$ (3.23)
$m-1-r$ $=\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}.$ (3.24)
These are polynomial equations of degree $2$.
Proof.
The “Four-Square Theorem” of Lagrange [1770] allows us to say:
Every $n\in\mathbb{N}$ is a sum of four squares.
Constraint (4.48) forces $r\geq 0$. Constraint (4.49) forces $r\leq m-1$, hence $r<m$. Given $r^{*}\in\{0,\ldots,m-1\}$, the four-square decompositions can be computed in polynomial time. By the algorithm of Rabin and Shallit [1986] we decompose an integer $n\in\mathbb{N}$ into four squares
$n=a^{2}+b^{2}+c^{2}+e^{2}$ (3.25)
This ensures the encoding is effective. Computable witnesses exist for any valid bound, and can be constructed in polynomial time (randomized) or quasi-polynomial time (deterministic under ERH). ∎
Remark.
In the completeness proof of Theorem 4.16, the witness construction algorithm (Exposition 4.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$. By the above algorithm, this can be done in time $O(\log((i+1)d))$ per decomposition, yielding total time $O(N\cdot\log(\max_{i}f_{i}))$ for all $N$ proof lines. This is polynomial in the proof parameters, as required.
3.2 Unique Representation
Remark.
As a “rule of thumb”: A sum satisfies Zeckendorf representation (i.e., uses non-consecutive Fibonacci numbers) when there are “gaps” in the sequence of Fibonacci numbers used. e.g., the tuple $\{F_{6},F_{8},F_{10}\}=\{8,21,55\}$.
Definition 3.3 (Zeckendorf Representation).
Every positive integer $n$ admits a unique representation as a sum of non-consecutive Fibonacci numbers:
$n=\sum_{\kappa\in S}F_{\kappa},$ (3.26)
where $S\subseteq\{2,3,4,\ldots\}$ contains no consecutive integers (i.e., if $\kappa\in S$, then $\kappa+1\notin S$), and $F_{k}$ denotes the $k$-th Fibonacci number with $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$.
Lemma 3.4 (Polynomial Encoding of Zeckendorf Digits).
For each integer $n$ with Zeckendorf representation involving indices up to $K$, introduce binary variables $d_{2},\ldots,d_{K}\in\{0,1\}$ satisfying:
  1. 1.
    Boolean constraint: $d_{\kappa}-d_{\kappa}^{2}=0$ for all $\kappa\in\{2,\ldots,K\}$,
  2. 2.
    Non-adjacency: $d_{\kappa}\cdot d_{\kappa+1}=0$ for all $\kappa\in\{2,\ldots,K-1\}$,
  3. 3.
    Reconstruction: $n=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{\kappa}$.
These constraints are polynomial of degree at most $2$.
Proof.
The boolean constraint $d_{\kappa}(1-d_{\kappa})=0$ forces $d_{\kappa}\in\{0,1\}$ over $\mathbb{N}$. The product $d_{\kappa}d_{\kappa+1}=0$ enforces non-adjacency. The sum is linear in the $d_{\kappa}$ variables. ∎
Exposition.
The Zeckendorf representation provides an alternative to prime-power encodings for sequences. We use it to enable Fibonacci addition, which can be implemented via digit-wise operations respecting the non-adjacency constraint (see Section 4.3).
3.3 Combined Encoding
Exposition.
For a proof sequence $\langle f_{1},\ldots,f_{N}\rangle$ where each $f_{i}$ is a Gödel number:
  1. 1.
    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (4.47) and Lemma 3.2.
  2. 2.
    Represent each $f_{i}$ in Zeckendorf form using Lemma 3.4.
This dual encoding allows:
  1. 1.
    Variable-length proofs (via existential quantification over $N,c,d$),
  2. 2.
    Arithmetic verification (via Fibonacci addition on Zeckendorf digits).
Regarding the consistency of “Dual Encoding:” Each Gödel number $f_{i}$ is encoded twice:
  1. 1.
    Via the $\beta$-function: $\beta(c,d,i)=f_{i}$ (using variables $c,d,r_{i},q_{i},\ldots$),
  2. 2.
    Via Zeckendorf: $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}$ (using digit variables $d_{i,\kappa}$).
These encodings are compatible: for any sequence $\langle f_{1},\ldots,f_{N}\rangle$, witnesses $c^{*},d^{*}$ exist (by Theorem 3.1), and each $f_{i}^{*}$ has a unique Zeckendorf representation (by Definition 3.3). The polynomial system constrains a single variable $f_{i}$ to satisfy both encodings simultaneously, ensuring consistency.
3.4 Summary of Polynomial Constraints
Constraint Expanded Equation $\boldsymbol{\delta}$ $\mathbb{N}$-Mechanism
Pre guard-wrapping
$\beta$-Division $c-q_{i}(1+(i+1)d)-r_{i}=0$ 2 Linear in $c,r_{i}$; quadratic in $q_{i},d$
Lower bound $r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}=0$ 2 Sum of squares $\geq 0$
Upper bound $(i+1)d-1-r_{i}-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\delta_{i}^{2}=0$ 2 Sum of squares $\geq 0$
Boolean $d_{i,\kappa}-d_{i,\kappa}^{2}=0$ 2 Forces $d_{i,\kappa}\in\{0,1\}$
Non-adjacency $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ 2 Product vanishes
Zeckendorf sum $f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}=0$ 1 Linear combination
Axiom membership $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$ 3 Boolean times sum-of-squares
MP arithmetic $f_{i}-f_{j}-f_{k}=0$ 1 Linear equality
Exactly-one $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}-1=0$ 1 Linear in booleans
Target check $f_{N}-g_{T}=0$ 1 Linear equality
Post guard-wrapping
$\beta$-Division guard $u_{\beta,i}\cdot(c-q_{i}(1+(i+1)d)-r_{i})=0$ 3 $u_{\beta,i}=1+v_{\beta,i}^{2}\geq 1$
Lower-bound guard $u_{L,i}\cdot(r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2})=0$ 3 $u_{L,i}=1+v_{L,i}^{2}\geq 1$
Upper-bound guard $u_{U,i}\cdot((i+1)d-1-r_{i}-\alpha_{i}^{2}-\cdots)=0$ 3 $u_{U,i}=1+v_{U,i}^{2}\geq 1$
Boolean guard $u_{B,i,\kappa}\cdot(d_{i,\kappa}-d_{i,\kappa}^{2})=0$ 3 $u_{B,i,\kappa}=1+v_{B,i,\kappa}^{2}\geq 1$
Non-adjacency guard $u_{N,i,\kappa}\cdot(d_{i,\kappa}\cdot d_{i,\kappa+1})=0$ 3 $u_{N,i,\kappa}=1+v_{N,i,\kappa}^{2}\geq 1$
Zeckendorf guard $u_{Z,i}\cdot(f_{i}-\sum_{\kappa}F_{\kappa}d_{i,\kappa})=0$ 2 $u_{Z,i}=1+v_{Z,i}^{2}\geq 1$
Axiom guard (Already $\delta=3$ before wrapping) 3 Boolean activation
MP-sum guard $u_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0$ 2 $u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1$
Exactly-one guard $u_{J,i}\cdot(b_{ax,i}+\sum_{j,k}b_{mp,i,j,k}-1)=0$ 2 $u_{J,i}=1+v_{J,i}^{2}\geq 1$
Target guard $\mathrm{T}\cdot(f_{N}-g_{T})=0$ 2 $\mathrm{T}=1+U^{2}\geq 1$
Positivity constraints
All guards $u_{\bullet}-1-v_{\bullet}^{2}=0$ 2 Forces $u_{\bullet}\geq 1$
Global multiplier $\mathrm{T}-1-U^{2}=0$ 2 Forces $\mathrm{T}\geq 1$
Table 2: Complete degree bookkeeping. Base constraints have $\delta\leq 3$; guard-wrapping preserves $\delta\leq 3$ for all main constraints, with guard positivity at $\delta=2$.
3.5 Guard-Gadget Framework
Definition 3.5 (Boolean Constraint).
For any variable $b$ intended to represent a boolean value, the constraint
$b-b^{2}=0,\quad b\in\{0,1\}\subseteq\mathbb{N}$ (3.27)
Definition 3.6 (Guarded Constraint).
For any polynomial $E(\mathbf{x})$ of degree $\delta\leq 2$, introduce guard variable $u$ and slack variable $v\in\mathbb{N}$. The guarded constraint is the system:
$u\cdot E(\mathbf{x})$ $=0,$ (3.28)
$u-1-v^{2}$ $=0.$ (3.29)
Degree bound: If $\delta\leq 2$, then $\delta(u\cdot E)\leq 1+2=3$ and $\delta(u-1-v^{2})=2$.
Proposition 3.7 (Guard Semantics).
The guarded constraint enforces:
  1. 1.
    $u=1+v^{2}\geq 1$ for all solutions (from equation (3.29)),
  2. 2.
    $E(\mathbf{x})=0$ (forced by equation (3.28) since $u\geq 1$).
Proof.
Equation (3.29) forces $u=1+v^{2}\geq 1$. Since $u\neq 0$, equation (3.28) yields $E(\mathbf{x})=0$. ∎
Definition 3.8 (Global Multiplier).
To enable selective activation of guards, introduce a global multiplier $T\geq 1$ via:
$T=1+U^{2},\quad U\in\mathbb{N}.$ (3.30)
This ensures $T\geq 1$ for all solutions. Optionally, this can be enforced via guarded constraints:
$u_{T}\cdot(T-1-U^{2})$ $=0,$ (3.31)
$u_{T}-1-v_{T}^{2}$ $=0,$ (3.32)
where equation (3.32) forces $u_{T}=1+v_{T}^{2}\geq 1$, and equation (3.31) then enforces $T=1+U^{2}$.
Proposition 3.9 (Selective Activation).
Given a guard $u_{i}$ with $u_{i}=1+v_{i}^{2}$ (from Definition 3.6) and a global multiplier $T\geq 1$, the constraint
$T\cdot(1-u_{i})=0$ (3.33)
forces $u_{i}=1$ (equivalently, $v_{i}=0$).
Proof.
Since $T\geq 1$ and $T\cdot(1-u_{i})=0$, we have $1-u_{i}=0$, hence $u_{i}=1$. From $u_{i}=1+v_{i}^{2}$, this implies $v_{i}=0$. ∎
Exposition.
The selective activation mechanism is used to enforce that specific constraints hold in a solution. For example, in proof verification (Section 4.4), we use it to force exactly one justification (axiom or modus ponens) for each line of a proof. The guard-gadget framework provides a systematic method to enforce boolean constraints without degree inflation (Definition 3.5), ensure all polynomial constraints are non-negative and solvable over $\mathbb{N}$ (Proposition 3.7), and selectively activate constraints via a global multiplier (Proposition 3.9). All gadgets maintain degree $\leq 3$ when applied to linear or quadratic base constraints. This forms the technical foundation for encoding proof verification at cubic degree.
4 Construction
We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{T}$.
4.1 Proof Sequence Encoding
We encode a proof as a sequence $\langle f_{1},\ldots,f_{N}\rangle$ of Gödel numbers.
Lemma 4.1 ($\beta$-Function Encoding).
For each $i\in\{1,\ldots,N\}$, enforce:
$u_{\beta,i}\cdot(c-q_{i}(1+(i+1)d)-r_{i})$ $=0,$ (4.34)
$u_{\beta,i}-1-v_{\beta,i}^{2}$ $=0,$ (4.35)
[…]
Lemma 4.2 (Four-Squares Bounds).
For each $i\in\{1,\ldots,N\}$, enforce:
$u_{L,i}\cdot(r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2})$ $=0,$ (4.36)
$u_{L,i}-1-v_{L,i}^{2}$ $=0,$ (4.37)
$u_{U,i}\cdot((i+1)d-r_{i}-1-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\eta_{i}^{2})$ $=0,$ (4.38)
$u_{U,i}-1-v_{U,i}^{2}$ $=0.$ (4.39)
Degree analysis: For all equations: $\delta\leq 3$.
Lemma 4.3 (Zeckendorf Representation).
For each $i\in\{1,\ldots,N\}$ and $\kappa\in\{2,\ldots,K\}$, enforce:
$u_{B,i,\kappa}\cdot(d_{i,\kappa}-d_{i,\kappa}^{2})$ $=0,$ (4.40)
$u_{B,i,\kappa}-1-v_{B,i,\kappa}^{2}$ $=0,$ (4.41)
and for $\kappa\in\{2,\ldots,K-1\}$:
$u_{N,i,\kappa}\cdot(d_{i,\kappa}\cdot d_{i,\kappa+1})$ $=0,$ (4.42)
$u_{N,i,\kappa}-1-v_{N,i,\kappa}^{2}$ $=0.$ (4.43)
Finally, enforce the reconstruction:
$u_{Z,i}\cdot\left(f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}\right)$ $=0,$ (4.44)
$u_{Z,i}-1-v_{Z,i}^{2}$ $=0.$ (4.45)
Degree analysis: For all equations: $\delta\leq 3$.
Exposition.
Constraints (Lemmas 4.14.3) ensure that the sequence $\langle f_{1},\ldots,f_{N}\rangle$ is encoded via $\beta(c,d,i)=f_{i}$ (Theorem 3.1), and each $f_{i}$ has a valid Zeckendorf representation with digits $d_{i,\kappa}$ (Lemma 3.4). The dual encoding is consistent (Exposition). “Why Carry Propagation is Unnecessary.” The key insight is that Zeckendorf representation is unique (Definition 3.3). The constraints enforce $f_{i}=f_{j}+f_{k}$ (arithmetic sum, via Equation (4.59)), and $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}$ with $d_{i,\kappa}\in\{0,1\}$ and $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ (via Proposition 4.4). By uniqueness, there exists exactly one choice of $(d_{i,2},\ldots,d_{i,K})$ satisfying the second condition for the value $f_{i}$ determined by the first condition. Hence the constraints force $d_{i,\kappa}$ to be the Zeckendorf digits of $f_{j}+f_{k}$, without explicitly computing the addition “digit-by-digit”. A procedural approach would introduce auxiliary variables $s_{i,j,k,\kappa}$ (sum digits) and $c_{i,j,k,\kappa}$ (carries), computing $s_{\kappa}=d_{j,\kappa}+d_{k,\kappa}+c_{\kappa-1}$ and normalizing to Zeckendorf form. This requires degree-$4$ constraints to handle carry propagation. Our declarative approach avoids this by relying on existential quantification: the solver “guesses” digits $(d_{i,\kappa})$, and uniqueness ensures only the correct guess satisfies all constraints.
4.2 Combined Proof Encoding
We combine the $\beta$-function encoding (for storing the proof sequence) with Zeckendorf representation (for bounding formula sizes) to obtain a complete encoding of proof lines.
Proposition 4.4 (Combined Proof Line Encoding).
For each proof line $i\in\{1,\ldots,N\}$, the Gödel number $f_{i}\in\mathbb{N}$ satisfies:
$f_{i}=\beta(c,d,i)=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},$ (4.46)
where:
  1. 1.
    $c,d\in\mathbb{N}$ are the $\beta$-function parameters encoding the sequence $(f_{1},\ldots,f_{N})$,
  2. 2.
    $d_{i,\kappa}\in\{0,1\}$ are the Zeckendorf digits of $f_{i}$,
  3. 3.
    $K=\lceil\log_{\phi}(\max_{i}f_{i})\rceil+2$ bounds the number of Fibonacci digits needed.
The constraint system enforces:
$f_{i}-q_{i}\cdot(1+(i+1)\cdot d)-r_{i}$ $=0,$ (4.47)
$r_{i}-a_{i}^{2}-b_{i}^{2}-c_{i}^{2}-e_{i}^{2}$ $=0,$ (4.48)
$(i+1)\cdot d-r_{i}-1-\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}-\epsilon_{i}^{2}$ $=0,$ (4.49)
$d_{i,\kappa}\cdot(d_{i,\kappa}-1)$ $=0,$ (4.50)
$d_{i,\kappa}\cdot d_{i,\kappa+1}$ $=0,$ (4.51)
$f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}$ $=0.$ (4.52)
Degree analysis: All constraints have $\delta\leq 2$.
Proof.
Equations (4.47)–(4.49) encode $f_{i}=\beta(c,d,i)$ via the $\beta$-function. Equations (4.50)–(4.52) enforce the Zeckendorf representation of $f_{i}$. Uniqueness of Zeckendorf representation ensures these constraints are consistent. ∎
4.3 Axiom Membership
We verify that a proof line is an axiom by testing membership in the finite set of axiom Gödel numbers $\{g_{1},\ldots,g_{M}\}$.
Lemma 4.5 (Axiom Test Constraint).
For each line $i\in\{1,\ldots,N\}$ we introduce a Boolean activation variable $b_{ax,i}\in\{0,1\}$ and enforce:
$b_{ax,i}\cdot(b_{ax,i}-1)$ $=0,$ (4.53)
$b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}$ $=0.$ (4.54)
[…]
Proposition 4.6.
We probe for completeness and soundness.
  1. 1.
    The constraint system in Lemma 4.5 correctly encodes (sound) axiom membership if $b_{ax,i}=1$ in a solution, then:
    1. (a)
      Equation (4.54) forces $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$.
    2. (b)
      Since each term $(f_{i}-g_{\ell})^{2}\geq 0$, this holds if and only if $f_{i}=g_{\ell}$ for some $\ell\in\{1,\ldots,M\}$.
    3. (c)
      Therefore $f_{i}$ is an axiom.
  2. 2.
    If line $i$ is an complete in respect to Gödel number $f_{i}=g_{\ell}$ for some $\ell$:
    1. (a)
      Set $b_{ax,i}=1$.
    2. (b)
      Then $(f_{i}-g_{\ell})^{2}=0$ and $(f_{i}-g_{\ell^{\prime}})^{2}>0$ for all $\ell^{\prime}\neq\ell$.
    3. (c)
      The sum $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$, so equation (4.54) is satisfied.
Proof.
Immediate from the properties of squares over $\mathbb{N}$. ∎
Exposition.
The constraint $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0$ encodes the logical disjunction:
$b_{ax,i}=1\implies(f_{i}=g_{1})\lor(f_{i}=g_{2})\lor\cdots\lor(f_{i}=g_{M}).$ (4.55)
This is a standard technique in Diophantine encoding: to express $x$ equals one of $a_{1},\ldots,a_{n}$, we use:
$\sum_{j=1}^{n}(x-a_{j})^{2}=0.$ (4.56)
The key property is that a sum of non-negative terms vanishes if and only if each term vanishes. Over $\mathbb{N}$, this requires $x=a_{j}$ for some $j$. Degree analysis: The sum $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}$ has degree $2$ (each term is a square of a linear expression). Multiplying by the boolean $b_{ax,i}$ increases the degree to $3$.
4.4 Modus Ponens Verification
We verify modus ponens by directly enforcing the arithmetic constraint $f_{i}=f_{j}+f_{k}$ for the appropriate triple $(i,j,k)$. The Zeckendorf digit constraints ensure that the digits $d_{i,\kappa}$ represent this sum uniquely, without requiring explicit carry propagation.
Lemma 4.7 (Modus Ponens Constraint).
For each triple $(i,j,k)$ with $1\leq j,k<i\leq N$, introduce:
  1. 1.
    Boolean activation variable $b_{mp,i,j,k}\in\{0,1\}$,
  2. 2.
    Guard variables $u_{mp,i,j,k},v_{mp,i,j,k}\in\mathbb{N}$.
Enforce the constraints:
$b_{mp,i,j,k}\cdot(b_{mp,i,j,k}-1)$ $=0,$ (4.57)
$u_{mp,i,j,k}-1-v_{mp,i,j,k}^{2}$ $=0,$ (4.58)
$u_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})$ $=0.$ (4.59)
[…]
Lemma 4.8 (Correctness of Modus Ponens Encoding).
The constraint system in Lemma 4.7 correctly encodes modus ponens: Soundness. If $b_{mp,i,j,k}=1$ in a solution, then, equation (4.58) necessitates
$u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1;$ (4.60)
equation (4.59) with $u_{mp,i,j,k}\geq 1$ forces $f_{i}=f_{j}+f_{k}$, while proposition 4.4, the digits $d_{i,\kappa}$ satisfy:
$f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (4.61)
By uniqueness of Zeckendorf representation, the digits $d_{i,\kappa}$ are the unique Zeckendorf digits of $f_{j}+f_{k}$. Completeness. Given a proof where line $i$ is derived from lines $j,k$ via modus ponens (so $f_{i}=f_{j}+f_{k}$ in $\mathbb{N}$): Set $b_{mp,i,j,k}=1$ and $b_{mp,i,j^{\prime},k^{\prime}}=0$ for all $(j^{\prime},k^{\prime})\neq(j,k)$. Compute the Zeckendorf representation of $f_{j}+f_{k}$:
$f_{j}+f_{k}=\sum_{\kappa\in S}F_{\kappa},\quad S\subseteq\{2,\ldots,K\},\;\kappa\in S\implies\kappa+1\notin S.$ (4.62)
We assign $d_{i,\kappa}=\mathbb{1}_{\kappa\in S}$ (indicator function). Set $u_{mp,i,j,k}=1$, $v_{mp,i,j,k}=0$. All constraints (4.57)–(4.59) are satisfied. See [Zeckendorf, 1972] for existence and uniqeness.
Proof.
By soundness and completeness. Steps (1)–(2) follow directly from the guard mechanism. Step (3) follows from the Zeckendorf constraints in Proposition 4.4, which apply to $f_{i}$ regardless of how $f_{i}$ is justified. Step (4) is immediate from the uniqueness theorem. By construction, $b_{mp,i,j,k}\cdot(b_{mp,i,j,k}-1)=1\cdot 0=0$. The guard equation gives $u_{mp,i,j,k}=1+0^{2}=1$. The sum constraint becomes:
$1\cdot(f_{i}-f_{j}-f_{k})=1\cdot\left(\sum_{\kappa\in S}F_{\kappa}-f_{j}-f_{k}\right)=1\cdot 0=0,$ (4.63)
where the second equality uses the definition of $S$ in step (2). ∎
Exposition.
Regarding the elimination of the “carry propagation”: A procedural approach to modus ponens would compute the Fibonacci sum $f_{i}=f_{j}+f_{k}$ digit-by-digit, introducing auxiliary variables $s_{i,j,k,\kappa}$ (sum digits) and $c_{i,j,k,\kappa}$ (carry bits) with constraints:
$s_{i,j,k,\kappa}$ $=d_{j,\kappa}+d_{k,\kappa}+c_{i,j,k,\kappa-1},$ (4.64)
$c_{i,j,k,\kappa}$ $=\begin{cases}1&\text{if }s_{i,j,k,\kappa}\geq 2,\\ 0&\text{otherwise}.\end{cases}$ (4.65)
Encoding the carry logic requires $\delta\geq 4$ to normalize adjacent $1$-s in the binary representation (since Fibonacci addition can produce carries propagating multiple positions). Our key observation: Carry propagation is unnecessary. The constraint $f_{i}=f_{j}+f_{k}$ (linear in $f_{i},f_{j},f_{k}$) combined with the Zeckendorf digit constraints (quadratic in $d_{i,\kappa}$) suffices. The existential quantification $\exists d_{i,\kappa}$ allows the constraint solver to guess the correct digits, which are then verified by:
  1. 1.
    The reconstruction constraint $f_{i}=\sum_{\kappa}F_{\kappa}\cdot d_{i,\kappa}$ (Equation 4.52),
  2. 2.
    The non-adjacency constraint $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ (Equation 4.51).
This is analogous to how Diophantine equations encode division: rather than computing $\lfloor x/y\rfloor$ procedurally, we introduce a quotient variable $q$ and enforce $x=qy+r$ with $0\leq r<y$ declaratively. The constraint system specifies what must hold, not how to compute it.
Benefit: The procedural approach requires degree-$4$ constraints. The declarative approach uses only:
  1. 1.
    Linear sum with guard: $u\cdot(f_{i}-f_{j}-f_{k})=0$ (degree $2$),
  2. 2.
    Quadratic digit constraints: already present in Proposition 4.4.
This maintains the overall degree bound of $3$ for the entire system.
Procedure 4.9 (Greedy Witness Construction).
The completeness direction of Lemma 4.8 relies on the computability of Zeckendorf representations. Given $n\in\mathbb{N}$, the Zeckendorf digits can be computed by the “greedy algorithm” [^5] An algorithm that makes locally optimal choices at each step, selecting the largest (or best) available option without reconsidering previous decisions. in $O(\log n)$ steps that produces the unique representation satisfying the non-adjacency constraint.:
[…]
Exposition.
For verification, given a formal proof $\pi$ of length $N$ with Gödel numbers $f_{1},\ldots,f_{N}$, we can effectively construct a solution $\vec{x}^{*}\in\mathbb{N}^{m}$ by computing $c^{*},d^{*}$ via the $\beta$-function, computing Zeckendorf digits $d_{i,\kappa}^{*}$ for each $f_{i}$ using the greedy algorithm, setting activation variables $b_{ax,i}^{*}$ or $b_{mp,i,j,k}^{*}$ according to the justification of line $i$, and computing all guard variables $u^{*},v^{*}$ to satisfy the positivity constraints. This witness construction is polynomial-time in $N$ and $\max_{i}f_{i}$, demonstrating that the reduction from provability to Diophantine solvability is effective.
4.5 Proof Justification
Each line of the proof must be justified either as an axiom or by modus ponens from earlier lines.
Lemma 4.10 (Modus Ponens Activation).
For each triple $(i,j,k)$ with $j,k<i$, introduce a boolean variable $b_{mp,i,j,k}\in\{0,1\}$ indicating whether line $i$ is derived from lines $j$ and $k$ via modus ponens. Enforce:
$u_{MP,i,j,k}\cdot(b_{mp,i,j,k}-b_{mp,i,j,k}^{2})$ $=0,$ (4.66)
$u_{MP,i,j,k}-1-v_{MP,i,j,k}^{2}$ $=0,$ (4.67)
and the MP test:
$b_{mp,i,j,k}\cdot(f_{i}-m_{i,j,k})$ $=0.$ (4.68)
[…]
Lemma 4.11 (Exactly-One Justification).
For each $i\in\{1,\ldots,N\}$, enforce:
$u_{J,i}\cdot\left(b_{ax,i}+\sum_{\begin{subarray}{c}j,k=1\\ j,k<i\end{subarray}}^{N}b_{mp,i,j,k}-1\right)$ $=0,$ (4.69)
$u_{J,i}-1-v_{J,i}^{2}$ $=0.$ (4.70)
[…]
Proposition 4.12 (Justification Correctness).
Constraint 4.11 necessitates that exactly one of the following holds for each line $i$:
  1. 1.
    $b_{ax,i}=1$ and $b_{mp,i,j,k}=0$ for all $j,k<i$ (line $i$ is an axiom), or
  2. 2.
    $b_{ax,i}=0$ and $b_{mp,i,j^{*},k^{*}}=1$ for exactly one pair $(j^{*},k^{*})$ (line $i$ is derived by MP).
Proof.
Since $b_{ax,i},b_{mp,i,j,k}\in\{0,1\}$ (by 4.10), equation (4.69) necessitates:
$b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1.$ (4.71)
This is a linear Diophantine equation with boolean variables, which has a solution if and only if exactly one variable is 11 and the rest are 0. ∎
Exposition.
We consider the activation mechanism. If $b_{ax,i}=1$, then Constraint 4.5 forces $f_{i}\in\mathcal{A}$. If $b_{mp,i,j,k}=1$, then Constraint 4.10 forces $f_{i}=m_{i,j,k}=f_{j}+f_{k}$ (assuming the Fibonacci addition constraints hold). Constraint 4.11 ensures that exactly one of these activation variables is $1$ for each line $i$. This encodes the informal proof rule: "Each line is either an axiom or follows from two earlier lines by MP."
4.6 Target Formula Constraint
Lemma 4.13 (Target Check).
Let $G_{T}$ be the target formula with Gödel number $g_{T}$. Enforce, with $T$ is the global multiplier (Definition 3.8),
$T\cdot(f_{N}-g_{T})$ $=0,$ (4.72)
$T-1-U^{2}$ $=0,$ (4.73)
[…]
Lemma 4.14 (Target Correctness).
Constraint 4.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
Proof.
Since $T=1+U^{2}\geq 1$ (by equation (4.73)), equation (4.72) forces $f_{N}=g_{T}$. ∎
4.7 Complete System Assembly
Definition 4.15 (Complete Constraint System).
Let $\Phi(\mathbf{x})$ denote the conjunction of all constraints from Sections 4.14.6:
[…]
The system $\Phi(\mathbf{x})=\mathbf{0}$ consists of polynomial equations, each of degree at most $3$.
Theorem 4.16 (Finite Cubic System for Proof Verification).
Let $T_{\text{Fib}}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be a sentence. There exists a finite system $\Phi(\mathbf{x})$ of polynomial equations over $\mathbb{N}$, each of degree at most $3$, such that:
$T_{\text{Fib}}\vdash G_{T}\quad\iff\quad\exists\mathbf{x}\in\mathbb{N}^{m}:\Phi(\mathbf{x})=\mathbf{0}.$ (4.75)
Proof.
By soundness and completeness. ($\Rightarrow$) Completeness:
  1. 1.
    Suppose $T_{\text{Fib}}\vdash G_{T}$. Then there exists a finite proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ where $\varphi_{N}=G_{T}$. Let $f_{i}$ be the Gödel number of $\varphi_{i}$. By Theorem 3.1, there exist $c^{*},d^{*}\in\mathbb{N}$ such that $\beta(c^{*},d^{*},i)=f_{i}$ for all $i\leq N$. For each $f_{i}$, compute its Zeckendorf representation $d^{*}_{i,\kappa}$ (which exists and is unique by Definition 3.3). For each line $i$:
    1. (a)
      If $\varphi_{i}\in\mathcal{A}$, set $b^{*}_{ax,i}=1$ and $b^{*}_{mp,i,j,k}=0$ for all $j,k$.
    2. (b)
      If $\varphi_{i}$ is derived by MP from $\varphi_{j}$ and $\varphi_{k}$, set $b^{*}_{ax,i}=0$, $b^{*}_{mp,i,j,k}=1$, and $b^{*}_{mp,i,j^{\prime},k^{\prime}}=0$ for $(j^{\prime},k^{\prime})\neq(j,k)$.
  2. 2.
    Compute the Fibonacci addition witnesses (carries $c^{*}_{i,j,k,\kappa}$, sums $s^{*}_{i,j,k,\kappa}$) Finally, compute all guard variables $u^{*}_{\bullet},v^{*}_{\bullet}$ to satisfy definitions 3.63.8. By construction, $\mathbf{x}^{*}=(N,c^{*},d^{*},\{f^{*}_{i}\},\ldots)$ satisfies
    $\Phi(\mathrm{x}^{*})=\mathbf{0}.$
  3. 3.
    ($\Leftarrow$) Soundness: Suppose $\mathbf{x}^{*}=(N^{*},c^{*},d^{*},\{f^{*}_{i}\},\ldots)$ satisfies $\Phi(\mathbf{x}^{*})=\mathbf{0}$. Then:
    1. (a)
      By Constraint 4.1, $\beta(c^{*},d^{*},i)=f^{*}_{i}$ for all $i\leq N^{*}$.
    2. (b)
      By Constraint 4.3, each $f^{*}_{i}$ has a valid Zeckendorf representation.
    3. (c)
      By Constraint 4.11, each line $i$ is justified either as an axiom (Constraint 4.5) or by MP (Constraint 4.10).
    4. (d)
      By Constraint 4.13, $f^{*}_{N^{*}}=g_{T}$.
  4. 4.
    Let $\varphi_{i}$ be the formula with Gödel number $f^{*}_{i}$. The sequence $\langle\varphi_{1},\ldots,\varphi_{N^{*}}\rangle$ is a valid proof of $G_{T}$ in $T_{\text{Fib}}$, since:
    1. (a)
      Each $\varphi_{i}$ is either an axiom (see 4.3) or derived from earlier formulas by MP (by 4.12),
    2. (b)
      $\varphi_{N^{*}}=G_{T}$ (by Lemma 4.14)
  5. 5.
    Therefore, $T_{\text{Fib}}\vdash G_{T}$.
Exposition.
The system $\Phi(\mathbf{x})$ contains:
  1. 1.
    Variables: $O(KN^{3})$ variables, dominated by the MP tuples
    $\{b_{mp,i,j,k},c_{i,j,k,\kappa},\ldots\};$
  2. 2.
    Equations: $O(KN^{3})$ equations, with the same dominating term.
For a fixed maximum proof length $N$ and Fibonacci bound $K=O(\log(\max_{i}f_{i}))$, the system size is polynomial in $N$.
5 Aggregation
We now reduce the finite system $\Phi(\mathbf{x})$ from Theorem 4.16 to a single cubic polynomial.
5.1 Merging via Sum-of-Squares
We reduce the system $\Phi(\mathbf{x})$ from Theorem 4.16 to a single equation via sum-of-squares.
Degree bound. By Table 2, all polynomials $P_{i}\in\Phi$ satisfy $\delta(P_{i})\leq 3$ after guard-wrapping (Definition 3.6).
Merger. The system $\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$ is equivalent to
$P_{\text{merged}}(\mathbf{x})=\sum_{i=1}^{m}P_{i}(\mathbf{x})^{2}=0,\quad\delta(P_{\text{merged}})\leq 6.$ (5.76)
Size. With $m=O(KN^{3})$ constraints (Remark Exposition) and $s_{i}=O(K)$ monomials per $P_{i}$, we obtain $O(K^{3}N^{3})$ total monomials.
5.2 Degree Reduction to Cubic
Proposition 5.1 (Shielding).
By the construction of Jones [1982] in “Universal Diophantine Equation”, for any polynomial $P(\mathbf{x})$ of degree $d\geq 4$, there exists $Q(\mathbf{x},\mathbf{y})$ of degree $\leq 3$ with $O(d\cdot s)$ auxiliary variables (where $s$ is the monomial count) such that:
$\exists\mathbf{x}:P(\mathbf{x})=0\quad\Longleftrightarrow\quad\exists(\mathbf{x},\mathbf{y}):Q(\mathbf{x},\mathbf{y})=0.$ (5.77)
Proof sketch.
For each monomial $\mathbf{x}^{\mathbf{a}}$ with $|\mathbf{a}|>3$, introduce $y$ via $y=\prod_{j\in S}x_{j}^{a_{j}}$ (where $|\mathbf{a}|_{S}\geq 2$), expressed as a degree-$2$ constraint. Iterate until all monomials have degree $\leq 3$. See [Jones, 1982] for details. ∎
Applying Theorem 5.1 to $P_{\text{merged}}$ (degree $6$, $O(KN^{9})$ monomials by squaring analysis) introduces $k=O(KN^{9})$ auxiliary variables.
5.3 Main Result
Theorem 5.2 (Cubic Encoding of Provability).
Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axioms, and $G_{T}$ its Gödel sentence. There exists a polynomial
$P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$
of degree $\leq 3$ such that:
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (5.78)
The polynomial has $m=O(KN^{6})$ variables and $O(K^{2}N^{12})$ monomials.
Proof.
By our additive approach, theorem 4.16 (finite cubic system), sum-of-squares merging via (5.76), and theorem 5.1. ∎
Corollary 5.3 (Cubic Incompleteness in $\mathbb{N}$).
The solvability problem for diophantines with $\delta=3$ over $\mathbb{N}$ is undecidable.
Proof.
Provability in consistent recursively axiomatizable theories is incomplete by Gödel [1931]. Theorem 5.2 reduces this from $\delta\leq 4$ to $\delta\leq 3$ completeness. ∎
Remark (Non-Uniformity).
The polynomial $P$ depends on the (unknown) proof length $N$. This is inherent: no algorithm can compute the “correct” $P$ from $T$ and $G_{T}$ alone, as this would solve the Halting Problem.
6 Computation
6.1 Degree Reduction Protocol
We reduce $P_{\text{merged}}$ from (5.76) to degree $\leq 3$ via systematic monomial shielding.
Definition 6.1 (Canonical Gadget).
For constraint $E(\mathbf{x})=0$ with $\delta(E)\leq 2$, define:
$Q_{E}=U\cdot E(\mathbf{x})-Z^{2}=0,\quad U,Z\in\mathbb{N}.$ (6.79)
Then $\delta(Q_{E})=\max(1+\delta(E),2)\leq 3$ and $Q_{E}=0\iff E=0$ when $U\geq 1$.
Procedure 6.2 (Monomial Shielding).
Reduces a single monomial $m(\mathbf{x})$ of degree $d>3$ to degree $\leq 3$ via recursive factorization, maintaining the invariant that every introduced constraint has degree $\leq 3$.
[…]
Procedure 6.3 (Polynomial Degree Reduction).
Reduces polynomial $R(\mathbf{x})$ with $\delta(R)>3$ to an equivalent system where all constraints have degree $\leq 3$.
[…]
6.2 Example
Example 6.4.
For $m=x^{3}y^{2}z$ (degree 6), Algorithm 6.2 proceeds:
  1. 1.
    Split: $p=x^{3}$ ($\delta=3$), $q=y^{2}z$ ($\delta=3$).
  2. 2.
    Shield $p=x^{3}$:
    • Sub-split: $x^{3}=x\cdot x^{2}$ ($\delta=1+2$)
    • Shield $x^{2}$: introduce $w_{1}$ with $w_{1}-x^{2}=0$ ($\delta=2$)
    • Form $w_{2}-x\cdot w_{1}=0$ ($\delta=3$)
    Result: $x^{3}\to w_{2}$ via $\{w_{1}-x^{2}=0,w_{2}-xw_{1}=0\}$.
  3. 3.
    Shield $q=y^{2}z$:
    • Sub-split: $y^{2}z=y\cdot yz$ ($\delta=1+2$)
    • Shield $yz$: introduce $w_{3}$ with $w_{3}-yz=0$ ($\delta=2$)
    • Form $w_{4}-y\cdot w_{3}=0$ ($\delta=3$)
    Result: $y^{2}z\to w_{4}$ via $\{w_{3}-yz=0,w_{4}-yw_{3}=0\}$.
  4. 4.
    Combine: $v-w_{2}\cdot w_{4}=0$ ($\delta=3$)
Final output: $m\to v$ with constraints
$\{w_{1}-x^{2}=0,\,w_{2}-xw_{1}=0,\,w_{3}-yz=0,\,w_{4}-yw_{3}=0,\,v-w_{2}w_{4}=0\}.$ (6.80)
6.3 Solvability and Complexity
Lemma 6.5 (Solvability Preservation).
Let $(\Phi,R^{\prime})=\textsc{ReduceDegree}(R)$. Then:
$\exists\mathbf{x}^{*}\in\mathbb{N}^{n}:R(\mathbf{x}^{*})=0\quad\Longleftrightarrow\quad\exists(\mathbf{x}^{*},\mathbf{y}^{*})\in\mathbb{N}^{n+\ell}:\Phi=\mathbf{0}\land R^{\prime}=0.$ (6.81)
Proof.
($\Rightarrow$) Set $y_{i}^{*}=(\text{product})_{i}(\mathbf{x}^{*})$ for each shield variable; constraints $y_{i}-(\text{product})_{i}=0$ are satisfied and $R^{\prime}=R$ by substitution.
($\Leftarrow$) Constraints force $y_{i}^{*}=(\text{product})_{i}(\mathbf{x}^{*})$; substituting into $R^{\prime}$ yields $R(\mathbf{x}^{*})=0$. ∎
Proposition 6.6 (Complexity).
For polynomial $R$ with $m$ monomials of degree $\leq d$, Algorithm 6.3 introduces $O(m\log d)$ auxiliary variables and constraints, each of degree $\leq 3$.
Proof.
Each monomial requires $O(\log d)$ shielding steps (degree halves per recursion). Total: $\sum_{i=1}^{m}O(\log d_{i})=O(m\log d)$. ∎
Remark (Application to $P_{\text{merged}}$).
For $P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ with $\delta(P_{i})\leq 3$ (from Table 2), we analyze the complexity of the polynomial reduction as follows. Each squared polynomial $P_{i}^{2}$ has degree at most $6$ and contains at most $s_{i}^{2}$ monomials, where $s_{i}$ denotes the monomial count of $P_{i}$. The total monomial count is therefore $m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2})$, where $\bar{s}=\max_{i}s_{i}=O(K)$ by the bounds on Zeckendorf digit sums. Since Section 4.7 establishes that $m_{0}=O(K^{3}N^{3})$ for the total number of constraints, Proposition 6.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$. Combining this with the $O(K^{3}N^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
Lemma 6.7 (Cubic Reduction).
There exists a Diophantine polynomial $P_{\text{hard}}(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ with:
  1. 1.
    $\delta(P_{\text{hard}})\leq 3$,
  2. 2.
    $m=O(K^{3}N^{3})$ variables,
  3. 3.
    $P_{\text{hard}}(\mathbf{x}^{*})=0$ has solution in $\mathbb{N}^{m}\iff M$ accepts within $T(N)$ steps.
Proof.
Apply Algorithm 6.3 to $P_{\text{merged}}$ from (5.76). Degree bound follows from Algorithm 6.2’s invariant; solvability equivalence from Proposition 6.5; variable count from Remark Remark. ∎
7 Result
Exposition.
We establish the central equivalence: the polynomial $P_{\text{hard}}$ constructed in Section 4 faithfully encodes provability in $T_{\text{Fib}}$. The proof proceeds bidirectionally through soundness (Section 7.1) and completeness (Section 7.2), yielding the main result (Theorem 5.2).
7.1 Soundness
Theorem 7.1 (Soundness).
If there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$, then $T_{\text{Fib}}\vdash G_{T}$.
Proof.
Assume $P_{\text{hard}}(\vec{x}^{*})=0$. We extract a valid proof sequence from the solution.
Step 1. Global multiplier. By Definition 3.8, the vanishing of $P_{\text{hard}}$ implies:
$T^{*}=1+U^{*2}\geq 1.$ (7.82)
Step 2. Proof length. Since $\mathrm{T}^{*}\geq 1$, Lemma 3.9 forces all guarded constraints to hold. By Lemma 4.1, the sequence $\langle f_{1}^{*},\ldots,f_{N}^{*}\rangle$ satisfies:
$\beta(c^{*},d^{*},i)=f_{i}^{*}\quad\text{for all }i\in\{1,\ldots,N^{*}\}.$ (7.83)
Step 3. Zeckendorf representability. By Lemma 4.3, each $f_{i}^{*}$ admits a unique representation:
$f_{i}^{*}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}^{*},$ (7.84)
where $d_{i,\kappa}^{*}\in\{0,1\}$ with $d_{i,\kappa}^{*}\cdot d_{i,\kappa+1}^{*}=0$ (non-adjacency constraint from Lemma 3.4).
Step 4. Justification. By Proposition 4.12, for each line $i\leq N^{*}$:
  1. 1.
    If $b_{ax,i}^{*}=1$: Lemma 4.6 yields $f_{i}^{*}\in\{g_{1},\ldots,g_{M}\}$ (axiom membership).
  2. 2.
    If $b_{ax,i}^{*}=0$: Lemma 4.11 forces exactly one triple $(j,k)$ with $b_{mp,i,j,k}^{*}=1$.
Step 5. Modus ponens. For each activated triple with $b_{mp,i,j,k}^{*}=1$, Lemma 4.8 enforces:
$f_{i}^{*}=f_{j}^{*}+f_{k}^{*},$ (7.85)
with Zeckendorf digits computed via Exposition Exposition.
Step 6. Target formula. By Lemma 4.14:
$f_{N^{*}}^{*}=g_{T}.$ (7.86)
Conclusion. The sequence $\langle f_{1}^{*},\ldots,f_{N^{*}}^{*}\rangle$ constitutes a valid proof of $G_{T}$ in $T_{\text{Fib}}$. ∎
7.2 Completeness
Theorem 7.2 (Completeness).
If $\mathrm{T}_{\text{Fib}}\vdash G_{T}$, then there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$.
Proof.
Let $\pi=\langle f_{1},\ldots,f_{N}\rangle$ be a proof of $G_{T}$. We construct $\vec{x}^{*}$ by explicit assignment.
Step 1. Global multiplier. Set $\mathrm{T}^{*}=1$, $U^{*}=0$ (satisfying Definition 3.8).
Step 2. Sequence encoding. By Theorem 3.1, there exist $c^{*},d^{*}\in\mathbb{N}$ such that:
$\beta(c^{*},d^{*},i)=f_{i}\quad\text{for all }i\in\{1,\ldots,N\}.$ (7.87)
Set $f_{i}^{*}=f_{i}$ for all $i\leq N$, and compute remainders $r_{i}^{*}$ and auxiliary variables via Lemma 4.2.
Step 3. Zeckendorf digits. For each $i\leq N$, compute the Zeckendorf representation using Procedure 4.9:
$d_{i,\kappa}^{*}=\mathbb{1}_{\kappa\in S_{i}},$ (7.88)
where $S_{i}\subseteq\{2,\ldots,K\}$ indexes the Fibonacci numbers in the unique decomposition of $f_{i}$.
Step 4. Justification variables. For each line $i$:
  1. 1.
    If $f_{i}$ is an axiom: set $b_{ax,i}^{*}=1$ and $b_{mp,i,j,k}^{*}=0$ for all $(j,k)$.
  2. 2.
    If $f_{i}$ is derived from lines $j,k$ via modus ponens: set $b_{ax,i}^{*}=0$, $b_{mp,i,j,k}^{*}=1$, and $b_{mp,i,j^{\prime},k^{\prime}}^{*}=0$ for $(j^{\prime},k^{\prime})\neq(j,k)$.
Step 5. Guard variables. For all guard constraints (Definitions 3.63.8), set:
$u_{\bullet}^{*}=1,\quad v_{\bullet}^{*}=0.$ (7.89)
Step 6. Verification. By construction:
  1. 1.
    Lemma 4.1 holds (via $\beta$-function assignment),
  2. 2.
    Lemma 4.3 holds (via Zeckendorf uniqueness, Definition 3.3),
  3. 3.
    Lemma 4.5 holds for axiom lines (via $b_{ax,i}^{*}=1$),
  4. 4.
    Lemma 4.7 holds for MP lines (via arithmetic constraint $f_{i}=f_{j}+f_{k}$, see Exposition),
  5. 5.
    Lemma 4.13 holds (since $f_{N}^{*}=g_{T}$).
Conclusion. All constraints vanish under $\vec{x}^{*}$, hence $P_{\text{hard}}(\vec{x}^{*})=0$. ∎
7.3 Main Algorithm: Encoding Gödel’s Theorem
We present the complete reduction in algorithmic form:
[…]
Theorem 7.3 (Correctness of Algorithm 7.3).
The polynomial $P_{\text{final}}$ output by the algorithm satisfies:
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P_{\text{final}}(\mathbf{x})=0.$ (7.90)
Proof.
By Theorem 4.16 (soundness and completeness of Stages 1–6), Equation (5.76) (sum-of-squares equivalence), and Proposition 6.5 (degree reduction preserves solvability). ∎
7.4 Undecidability Consequences
Corollary 7.4 (Undecidability of $\mathcal{D}_{3}$).
Let there be a class $\mathcal{D}_{3}$ of cubic Diophantine equations where $\delta=3$. For $\mathcal{D}_{3}$, there exists no algorithm deciding, for arbitrary $Q(\vec{x})\in\mathbb{Z}[\vec{x}]$ with $\delta(Q)\leq 3$, whether $Q(\vec{x})=0$ has a solution in $\mathbb{Z}^{n}$.
Proof.
By Theorems 7.1 and 7.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$. By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{T}$ is independent of $\mathrm{T}_{\text{Fib}}$. Since $\delta(P_{\text{hard}})\leq 3$ (Theorem 5.2), this exhibits a cubic equation whose solvability is undecidable. Encoding arbitrary Turing machine halting problems via similar constructions yields a uniform reduction from the halting problem to $\mathcal{D}_{3}$ solvability. ∎
Theorem 7.5 (Uniform Halting Reduction).
There exists a computable function $\Phi:\mathbb{N}\to\mathbb{Z}[\vec{x}]$ such that for any Turing machine $M$ with Gödel number $\ulcorner M\urcorner$:
$\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace M\text{ halts on empty input}\textquotedblright\ltxml@oqmark@close{}\quad\Longleftrightarrow\quad\exists\vec{x}\in\mathbb{Z}^{k}:\Phi(\ulcorner M\urcorner)(\vec{x})=0,$ (7.91)
where $\delta(\Phi(\ulcorner M\urcorner))\leq 3$ and $k$ is independent of $M$.
Proof sketch.
By [Matiyasevich, 1970], there exists a polynomial $P_{\text{MRDP}}(m,\vec{y})$ encoding Turing machine halting. Applying the guard-gadget framework (Section 3.5) and degree reduction (Theorem 5.1) yields a cubic polynomial $\Phi(\ulcorner M\urcorner)$ preserving solvability. The map $M\mapsto\Phi(\ulcorner M\urcorner)$ is computable by effective construction of guards and auxiliary variables. ∎
7.5 Impredicativity
Thesis 7.6 (Impredicativity of Degree Threshold).
Let $\mathrm{T}$ be any consistent, recursively axiomatizable theory extending Robinson arithmetic $\mathsf{Q}$. For each $k\in\mathbb{N}$, let $\mathcal{D}_{k}$ denote the class of Diophantine equations with $\delta\leq k$. Define:
[…]
Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $\mathrm{T}$ proves:
  1. 1.
    $\exists!k\;B(k)$ (uniqueness of threshold),
  2. 2.
    $\forall m<k\;\mathsf{Dec}(m)$ (decidability below threshold),
  3. 3.
    $\forall m\geq k\;\neg\mathsf{Dec}(m)$ (undecidability at and above threshold).
Proof sketch.
Suppose such $B(k)$ exists with unique $k_{0}$ satisfying $\mathrm{T}\vdash B(k_{0})$. By the diagonal lemma, construct a Diophantine equation $E_{k_{0}}$ with $\delta(E_{k_{0}})=k_{0}$ whose solvability encodes $\neg B(k_{0})$ within $\mathrm{T}$. If $E_{k_{0}}$ is solvable, then $\mathrm{T}\vdash\neg B(k_{0})$, contradicting uniqueness. If $E_{k_{0}}$ is unsolvable, then $\mathcal{D}_{k_{0}}$ contains a decidable equation, contradicting clause (3). Hence no such $B(k)$ is definable in $\mathrm{T}$. ∎
Exposition.
The impredicativity arises from self-reference: defining $k$ as “the first undecidable degree” enables constructing an equation of degree $k$ that encodes the negation of this very claim. This mirrors the Grelling–Nelson paradox and reflects the predicativity restriction in the type theory of Martin-Löf [1980], where universes $\mathcal{U}_{i}$ cannot quantify over themselves.
Thesis 7.7 (Diophantine Impredicativity).
In the Calculus of Inductive Constructions, attempting to form:
$\mathsf{Threshold}:\Sigma_{k:\mathbb{N}}\left(\Pi_{m<k}\mathsf{Decidable}(\mathcal{D}_{m})\right)\times\left(\Pi_{m\geq k}\neg\mathsf{Decidable}(\mathcal{D}_{m})\right)$ (7.93)
requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined. By Thesis 7.6, no closed term of this type exists in any consistent extension of $\mathsf{Q}$. The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
$E_{k_{0}}:\mathcal{D}_{k_{0}}\quad\text{with}\quad\mathsf{Solvable}(E_{k_{0}})\leftrightarrow\neg B(k_{0}),$ (7.94)
precisely the impredicative loop forbidden in stratified type theories.
Corollary 7.8 (Stratification Necessity).
Any formal system capable of deciding $\mathcal{D}_{k}$ for some $k$ must either:
  1. 1.
    Externalize the threshold: Define $B(k)$ in a strictly stronger metatheory (e.g., $\mathsf{PA}$ proves undecidability of $\mathcal{D}_{3}$ using Gödel sentences not formalizable in $\mathsf{Q}$), or
  2. 2.
    Adopt impredicative axioms: E.g., $\mathsf{ZFC}$ proves existence of a least undecidable degree by quantifying over all sets of equations via the power set axiom (which is impredicative).
The cubic bound $k=3$ established here is not an internal theorem of any predicative arithmetic.
Exposition.
Under the Brouwer-Heyting-Kolmogorov interpretation, a proof of $\mathsf{Dec}(k)\lor\neg\mathsf{Dec}(k)$ requires either:
  1. 1.
    A witness algorithm $M_{k}$ deciding all equations in $\mathcal{D}_{k}$, or
  2. 2.
    A refutation proving no such $M_{k}$ exists (via diagonalization).
At the threshold $k=3$, neither is constructively available:
  1. 1.
    No algorithm exists (by MRDP and Corollary 7.4),
  2. 2.
    No uniform refutation exists (by Thesis 7.6, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
Thus, the threshold constitutes a Brouwerian counterexample to the law of excluded middle: a statement $P$ such that neither $P$ nor $\neg P$ admits constructive proof. The degree-$3$ bound is visible only from a classical metatheory (e.g., $\mathsf{ZFC}$ proves $\mathcal{D}_{3}$ is undecidable” by encoding the halting problem), but no predicative subsystem of $\mathsf{PA}$ can internalize this fact.
Thesis 7.9 (Impredicativity of Degree Threshold).
Let $\mathrm{T}$ be any consistent, effectively axiomatizable theory extending Robinson arithmetic $\mathsf{Q}$. For each $k\in\mathbb{N}$, let $\mathcal{D}_{k}$ denote the class of Diophantine equations of total degree at most $k$. Define the predicate:
$\text{\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace$M$ decides solvability for all equations in $\mathcal{D}_{k}$\textquotedblright\ltxml@oqmark@close{}}\iff\mathsf{Dec}(k)$
Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $\mathrm{T}$ proves:
  1. 1.
    $\exists!k\;B(k)$ (uniqueness of threshold),
  2. 2.
    $\forall m<k\;\mathsf{Dec}(m)$ (decidability below threshold),
  3. 3.
    $\forall m\geq k\;\neg\mathsf{Dec}(m)$ (undecidability at and above threshold).
Exposition.
This result formalizes the intuition that the boundary of undecidability for Diophantine equations is inherently impredicative. Any attempt to define the “first” degree $k$ at which undecidability arises leads to a form of self-reference: specifying such a $k$ enables the construction of a Diophantine equation of degree $k$ whose solvability encodes the negation of that very definition. Consequently, the threshold of undecidability is not first-order definable within any sufficiently strong theory. This mirrors classical semantic paradoxes—such as the Grelling–Nelson paradox [^6] Consult [Quine, 1976] for a thorough discussion of paradoxes.—in which the act of defining:
Is the adjective ’heterological’ a heterological word?
yields contradiction. Similarly, “Does undecidability start at kk?” becomes undecidable when answered affirmatively. If $k$ is the first undecidable degree, the evaluation criterion shifts, making the boundary undecidable if and only if it is decidable—a fixed point in arithmetic. Under HBK [^7] The Heyting–Brouwer–Kolmogorov (HBK), formalized by Heyting [1930], omits proof by the law of excluded middle unless constructively justified., this reflects the failure by appealing to the law of excluded middle (LEM): asserting $\mathsf{Dec}(k)\vee\neg\mathsf{Dec}(k)$ requires a uniform method for deciding solvability across all equations in $\mathcal{D}_{k}$, or uniformly refuting such a method—a totality that can be rejected without explicit witnesses.
Thus, impredicativity arises not merely from syntactic self-reference, but from the semantic attempt to define a global threshold for undecidability. The paradox is resolved constructively by recognizing that no such threshold is constructively definable: the transition from decidable to undecidable is not marked by a computable boundary, but by a failure of totality. In this sense, the undecidability threshold is a meta-mathematical horizon—a boundary that cannot be crossed without stepping outside the constructive framework. Any attempt to define it collapses into a fixed point of the form:
At what point does the undecidability of undecidability begin?